OpenStudy (anonymous):

x+6/12 divided by x-8/10 ???

OpenStudy (mertsj):

Do you know that dividing is the same as multiplying by the reciprocal of the divisor?

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow {x + 6 \over 12} \div {x - 8 \over 10} = {x + 6 \over 12} \times {10 \over x - 8} }\) This is a property.

OpenStudy (anonymous):

Yeah

OpenStudy (goformit100):

(5x+6)/(6x-46)

OpenStudy (goformit100):

(5x+6)/(6x-48) Ans.

OpenStudy (anonymous):

Thanks got it !

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow {x + 6 \over6} \times {5 \over x-8} }\) \(\Large \color{MidnightBlue}{\Rightarrow {5x + 30 \over 6x - 48} }\) @goformit100 WRONG

Parth (parthkohli):

You got the numerator wrong @goformit100

OpenStudy (goformit100):

(5x+48)/(6x-48) Ans.

OpenStudy (anonymous):

whats the numerator then?

Parth (parthkohli):

Again wrong

Parth (parthkohli):

5x + 30 is the numerator

OpenStudy (goformit100):

(5x+30)/(6x-48) Ans.

OpenStudy (goformit100):

now its correct, i s am sure

razor99 (razor99):

u stiil confusing

razor99 (razor99):

hello @Jshakour

razor99 (razor99):

so the anwer is